wire 42 ft long is divided into 2 sections. one section is 2/5 of the other. what are the lengths of the 2 sections?
(Algebra 2 question)

Answer by ewatrrr(24785)


Hi,
divided into 2 sections. one section is 2/5 of the other
Let x and (2/5)x represent the 2 sections
x + (2/5)x = 42ft
Mutilpying thru with 5 to clear the demoninator
5x + 2x = 5*42ft
7x = 210fy
x = 30ft
Other sections is (2/5)30 = 12ft
checking our answer
30ft + 12ft = 42ft